# Tracking Protostellar Variability in Massive Protoclusters with ALMA: I. Insights from QUARKS and MaMMOtH

### v1

@@ 3.1 Spectral Line Flagging and Self-calibration
The data were initially calibrated using the ALMA pipeline within the Common Astronomy Software Applications package (CASA, version 6.5.4.9; 2022PASP..134k4501C). This process included bandpass, gain, and flux calibration, employing the calibrators listed in Table A. The output of this procedure consists of calibrated measurement sets for each individual execution blocks. Each calibrated measurement set was subsequently split into individual files corresponding to the science targets, using the four scientific SPWs (SPW25, SPW27, SPW29, and SPW31). For each source, the four SPWs were combined to improve the continuum sensitivity.
@@
Line emission frequency ranges were flagged to clearly separate continuum emission from spectral line features. To ensure consistency across multiple epochs and surveys, a uniform flagging strategy was adopted for the spectral line frequency ranges. Initially, all prominent spectral line transitions within the four SPWs were identified in a reference epoch. This identification was performed by comparing the reduced data cubes with laboratory spectral databases (CDMS; 2001A&A...370L..49M), following the procedure outlined in Liu2024QUARKS-I. We manually adjusted the flagged ranges to include additional line emission present in other epochs but absent in the reference. This procedure was repeated iteratively until all visible spectral line features were consistently flagged across all epochs.
@@
Self-calibration is a technique used to correct visibility phases and/or amplitudes by comparing the observed visibilities with a model of the source itself (e.g., selfcal). Following standard interferometric guidelines (e.g., 1999ASPC..180.....T), phase-only self-calibration is applicable when the target is detected with a signal-to-noise ratio (S/N) ¿ 3 within a solution interval shorter than the timescale of significant phase variation across all baselines to a single antenna. Given the brightness of our targets, self-calibration was applied to the full dataset to enhance image quality and dynamic range. Two rounds of phase-only self-calibration were performed: the first with solint = “inf” and the second with solint = “int”. After each iteration, deeper cleaning was performed to refine the source model.
3.2 Continuum Imaging
@@
To minimize discrepancies in uv-plane sampling arising from differences in ALMA configurations and the Earth’s rotation among epochs, and to ensure reliable flux comparisons, we constrained the uvrange parameter during imaging within the tclean task. Based on the Amplitude v.s. UVWave distributions of each epoch, we identified and applied a common overlapping uv-range specific to each source. The same uv-range was used across epochs for a given source, but it could differ between sources. This procedure matches the baseline coverage in the uv-plane, thereby improving the cross-epoch consistency of the images. Our method provides a simple solution for the current data. Previous studies (e.g., 2019Francis; 2020Francis; 2022Francis) suggest that a more careful consideration of the uv-plane would be necessary in certain contexts, such as when the differences in array configuration between epochs are significant.
@@
Initial continuum images for each epoch were generated using the tclean task in CASA 6.6.1.17 with briggs robust weighting of 0.5. The Multi-scale Multi-Frequency Synthesis (mtmfs) deconvolution algorithm was adopted with nterm of 2. This algorithm improves wideband imaging by simultaneously modeling both spectral and spatial structures, restoring extended emission, and enhancing overall image fidelity. During the cleaning process, masks were automatically generated using the auto-multithresh algorithm, with input parameters recommended by the guides111https://casaguides.nrao.edu/index.php/Automasking_Guide_CASA_6.6.1 for the 12-m array (C-5 configuration). The image size was set to [900, 900] pixels with a pixel scale of 0.05″, and the pblimit parameter was set to 0.2.
@@
4 Results and Analysis
4.1 Source Extraction Strategy
@@
To ensure consistent and reliable comparisons of continuum emission across epochs, we developed a unified source extraction strategy that includes image alignment, structure identification, and flux measurement. Figure 2 presents the after primary beam correction continuum images of the source I15520–5234 at four epochs. This target, observed over a span of more than two years, is among the most comprehensively sampled sources in our study and serves as a representative example of the source extraction. Our first step is to align all continuum images for each target so that their spatial referencing is consistent across epochs, which is necessary because slight pointing variations exist among different observations. The alignment procedure was as follows. We visually inspected the continuum images and identified the morphological center of the source within the field of view using the CARTA image viewer333CARTA: Cube Analysis and Rendering Tool for Astronomy, https://cartavis.org/. Taking the first epoch as the reference, we defined a circular region centered at this position with an initial radius equal to 0.5 times the primary beam FWHM (13.24). We then verified that this region encompassed the entire source structure in all epochs. If not, the radius was incrementally increased until full coverage was achieved. In this case, the initial radius proved sufficient, as indicated by the cyan circle in Figure 2. Finally, we applied the Cutout2D package to crop all images to this common region, thereby ensuring spatial consistency across epochs and allowing accurate image alignment.
@@
Figure 3 illustrates our source extraction procedure, taking I15520–5234 as an example. To avoid elevated edge noise introduced by primary beam correction, structure identification was performed on the uncorrected continuum images. Structures were extracted independently for the MaMMOtH and QUARKS datasets. For each survey, multiple observations of the same source were combined to improve sensitivity. We used the astrodendro 444http://www.dendrograms.org/ algorithm to decompose the emission hierarchically. The highest hierarchical level is a “leaf” (i.e., a structure with no substructure), corresponding to what we define as a condensation. The three key parameters are min_value (the minimum pixel intensity to be considered), min_delta (the minimum height for any local maximum to be defined as an independent entity), and min_npix (the minimum number of pixels for a leaf to be defined as an independent entity). To ensure consistency across the two surveys, we adopted a uniform set of parameters: min_value = , min_delta = –, and min_npix equals to the pixel number of the beam area. Here, denotes the rms noise level of the continuum image. For sources with multiple epochs, the images were combined before structure identification, and was measured from the combined image. In the example shown in Figure 3, we used min_delta = , with for MaMMOtH and for QUARKS, and min_npix = 119.
@@
In Figure 3, structures extracted from the MaMMOtH and QUARKS datasets are shown with red and green contours. Although the condensations detected in the two surveys do not always match, all structures from both datasets are retained in the final union mask, shown in yellow. We constructed this final mask by taking the union of the individual masks of identified regions in different epochs. This union mask was then applied to each epoch for flux extraction. To ensure that the measured fluxes represent the true emission distribution, this step was performed after primary beam correction. We identified 383 condensations within the 22 protoclusters. The measured peak intensities and coordinates of all condensations are summarized in Appendix B, where the source name indicates the region (i.e., massive protocluster) in which each condensation resides. The complete table is available online. For each protocluster, condensation IDs are assigned starting from 1 and are ordered according to their spatial distribution in the map images, following a right-to-left and bottom-to-top sequence.
4.2 Relative Calibration
@@
After aligning the observations and extracting the sources, we derived and applied a relative flux calibration factor to each dataset in order to accurately track the peak intensity variations of a given object across all epochs. This approach allows for more robust measurements of intrinsic variability within each field. Our procedure generally follows the method outlined by (2017Mairs). The relative flux calibration procedure consists of the following four steps:
@@
We began by dividing the 22 protoclusters into six groups based on their combinations of observation dates to ensure consistent temporal sampling and enable coherent relative flux calibration. To maintain consistency in calibration across epochs, we ensured that each group shared an identical set of calibrator sources.
@@
In each group, we selected sources that exhibited S/N greater than 30 in all available epochs. We then calculated the mean peak intensity over the valid epochs and normalized the peak intensity of each epoch by this mean. We choose a threshold of 0.1 for the standard deviation of the normalized intensities as a compromise between retaining more calibrator candidates and ensuring the reliability of the calibration. Sources with standard deviations below this threshold were classified as stable and used as relative flux calibrators.
@@
For each group, we identified a subset of stable calibrator sources as described in the previous section. For every stable calibrator source and epoch , we performed a mean flux normalized ratio defined as:
(1) where is the observed peak intensity of the source in epoch , and is the mean peak intensity of that source across all valid epochs in the group. This normalization mitigates the impact of absolute calibration uncertainty and highlights the relative flux variation across epochs. The flux calibration factor for each epoch was then obtained by averaging the mean flux normalized ratios across all stable calibrators in the group:
(2) This approach ensures that each epoch is scaled consistently based on the collective behavior of stable sources within the group.
@@
4.3 Searching for (Sub)Millimeter Variables
Investigating flux variability in our survey is confronted by several key challenges. Each ALMA target was observed at only a small number of epochs (2–4 per protocluster), and the background rms varies across different protoclusters. To address these limitations, we employed two complementary methods to cross-validate variability among our samples (383 condensations). First, we adopted the methodology outlined by 2018Johnstone, originally developed for the JCMT Transient Survey. Although that survey benefits from more extensive temporal coverage and highly uniform observing conditions, which result in greater statistical robustness, our use of consistent data reduction and relative flux calibration still allows us to adopt the same framework as a practical indicator for identifying candidate variable sources. Second, to further increase the reliability of our results, we applied a difference map analysis to independently verify variability in the identified candidates.
4.3.1 Standard Deviation Analysis
For each detected condensation, we extract the peak intensity measured in each epoch and compute two key statistical quantities: the mean peak intensity, and the standard deviation (). The uncertainty in measuring the peak intensity of these sources is primarily influenced by two factors: the background rms for faint condensations, relative calibration accuracy for bright condensations. We defined the fiducial standard deviation for each source, , which represents the expected measurement uncertainty in peak intensity, as:
where denotes the background rms associated with the source . The term represents the expected relative flux calibration uncertainty, which we adopt to be 0.5% (see Section 4.2 for details), and is the mean peak flux density of the source.
Figure 5 shows the normalized standard deviation () as a function of the mean peak intensity for each condensation. In this work, we adopt a conservative variability threshold of , beyond which deviations are unlikely to result from calibration uncertainties or stochastic noise. The JCMT Transient Survey employed a lower threshold of 2 (2018Johnstone; 2025Chen); however, given our smaller number of epochs, we adopted a more stringent criterion to minimize false positives and ensure that only the most statistically significant outliers are flagged as candidate variables.
| Source Name | Condensation IDa | SD/SDfid | Fractional Difference (%)b | Mean Peak Intensity | Fractional Amplitude (%) | Typec | ||
| Epoch 21 | Epoch 31 | Epoch 41 | (mJy beam-1) | |||||
| I11332-6258 | 7 | 7.4 | 11.7 | 14.4 | 25.55 | 13.3 | Hot core | |
| I12320-6122 | 5 | 6.5 | 16.2 | 17.97 | 15.0 | Hot core | ||
| I13111-6228 | 10 | 28.7 | 68.3 | 19.51 | 51.1 | HC Hii region | ||
| I16076-5134 | 6 | 5.7 | -0.8 | 7.8 | 5.3 | 65.33 | 8.3 | Hot core |
| I16484-4603 | 1 | 7.4 | -7.3 | -9.3 | -10.4 | 97.37 | 11.1 | Hot core |
@@ 4.3.1 Standard Deviation Analysis
Applying this criterion, we identify five candidate variables among the 383 detections with multi-epoch measurements. Their quantitative variability metrics are summarized in Table 1. The fractional difference is calculated as , where and are the peak intensities at two different epochs. Multiple pairs of epochs (e.g., 21, 31, and 41) are used to evaluate the variability across the monitoring period. The observed fractional amplitude is defined as , where and correspond to the brightest and faintest peak intensities measured for each source, and denotes the mean peak intensity.
Among them, condensation 10 in I13111–6228, located at a distance of 2.97 kpc (Liu2024QUARKS-I), is particularly notable, exhibiting a normalized standard deviation of , far above the adopted threshold. This is the first detection of millimeter variability toward this protocluster. The other four candidates show more moderate excess deviations, with ranging from 5.7 to 7.4.
Although the fiducial standard deviation analysis provides a global identification of candidate variables, it does not explicitly capture the epoch-to-epoch flux behavior of individual condensations. To complement this approach, we extend the same fiducial standard deviation model by propagating uncertainties and computing the peak intensity ratio between epochs, referenced to epoch 1 for consistency (see Appendix C). We find that the variable candidates identified with this ratio-based method are fully consistent with those selected from the fiducial standard deviation analysis.
As an illustration, Figure 6 shows the source I13111–6228, in which condensation 10 exhibits the most significant intensity variation among all candidates. This condensation is clearly offset from the fiducial model, well exceeding the threshold, and is marked with a red star. For reference, the points with are colored green, reflecting their larger noise-dominated uncertainties, while those with higher S/N are shown in blue and provide more reliable measurements. Additional ratio plots for representative candidates are shown in Appendix C, and the complete set of figures is available online.
4.3.2 Difference Maps Analysis

### v3

@@ III.1 Spectral Line Flagging and Self-calibration
The data were initially calibrated using the ALMA pipeline within the Common Astronomy Software Applications package (CASA, version 6.5.4.9; CASA Team et al. 2022). This process included bandpass, gain, and flux calibration, employing the calibrators listed in Table A. The output of this procedure consists of calibrated measurement sets for each individual execution blocks. Each calibrated measurement set was subsequently split into individual files corresponding to the science targets, using the four scientific SPWs. For each source, the four SPWs were combined to improve the continuum sensitivity.
@@
Line emission frequency ranges were flagged to clearly separate continuum emission from spectral line features. To ensure consistency across multiple epochs and surveys, a uniform flagging strategy was adopted for the spectral line frequency ranges. Specifically, prominent spectral line features within the four SPWs were first identified in a reference epoch based on visual inspection of the visibility data. The flagged frequency ranges were then manually adjusted to include additional line emission present in other epochs but absent in the reference epoch. This procedure was repeated iteratively until all visible spectral line features were consistently flagged across all epochs.
@@
Self-calibration is a technique used to correct visibility phases and/or amplitudes by comparing the observed visibilities with a model of the source itself (e.g., Richards et al. 2022). Following standard interferometric guidelines (e.g., Taylor et al. 1999), phase-only self-calibration is applicable when the target is detected with a signal-to-noise ratio (S/N) ¿ 3 within a solution interval shorter than the timescale of significant phase variation across all baselines to a single antenna. Given the brightness of our targets, self-calibration was applied to the full dataset to enhance image quality and dynamic range. Two rounds of phase-only self-calibration were performed: the first with solint = “inf” and the second with solint = “int”. After each iteration, deeper cleaning was performed to refine the source model.
III.2 Continuum Imaging
@@
To minimize discrepancies in uv-plane sampling arising from differences in ALMA configurations and the Earth’s rotation among epochs, and to ensure reliable flux comparisons, we constrained the uvrange parameter during imaging within the tclean task. Based on the Amplitude v.s. UVWave distributions of each epoch, we determined the common overlapping uv-range for each source by selecting the maximum shared uv coverage across all available epochs. The same uv-range was used across epochs for a given source, but it could differ between sources. This procedure matches the baseline coverage in the uv-plane, thereby improving the cross-epoch consistency of the images. Our method provides a simple solution for the current data. Previous studies (e.g., Francis et al., 2019, 2020, 2022) suggest that a more careful consideration of the uv-plane would be necessary in certain contexts, such as when the differences in array configuration between epochs are significant.
@@
Initial continuum images for each epoch were generated using the tclean task in CASA 6.6.1.17 with briggs robust weighting of 0.5. The Multi-scale Multi-Frequency Synthesis (mtmfs) deconvolution algorithm was adopted with nterm of 2. This algorithm improves wideband imaging by simultaneously modeling both spectral and spatial structures, restoring extended emission, and enhancing overall image fidelity. During the cleaning process, masks were automatically generated using the auto-multithresh algorithm, with input parameters recommended by the guides111https://casaguides.nrao.edu/index.php/Automasking_Guide_CASA_6.6.1 for the 12-m array (C-5). The image size was set to [900, 900] pixels with a pixel scale of 0.05″, and the pblimit parameter was set to 0.2.
@@
IV Results and Analysis
IV.1 Source Extraction Strategy
To ensure consistent and reliable comparisons of continuum emission across epochs, we developed a unified source extraction strategy that includes image alignment, structure identification, and flux measurement. Figure 2 presents the after primary beam correction continuum images of the source I15520–5234 at four epochs. This target, observed over a span of more than two years, is among the most comprehensively sampled sources in our study and serves as a representative example of the source extraction.
@@ IV.1 Source Extraction Strategy
As a first step, we visually inspected the continuum images using the CARTA image viewer333CARTA: Cube Analysis and Rendering Tool for Astronomy, https://cartavis.org/. This visual inspection was used only to account for differences in the phase center settings between the QUARKS and MaMMOtH surveys and to define a common cutout region that encompasses all major source structures across epochs, rather than for the actual image alignment. Taking the first epoch as a reference, we identified the morphological center of the source and defined a circular region centered at this position with an initial radius equal to 0.5 times the primary beam FWHM (13.24). We verified that this region fully enclosed the source emission in all epochs; if not, the radius was incrementally increased until full coverage was achieved. In this case, the initial radius was sufficient, as indicated by the cyan circle in Figure 2. All images were then cropped to this common region using the Cutout2D444https://docs.astropy.org/en/stable/api/astropy.nddata.Cutout2D.html package, ensuring a consistent field of view across epochs.
During the construction of difference maps (Sect. IV.3.2), systematic relative offsets between epochs were identified, which manifested as positive–negative residual patterns (see Figure B1). For observations within the same survey, the offsets are more plausibly associated with variations in phase calibration quality between epochs. This motivated the use of a quantitative image alignment procedure. Image alignment was performed using a sub-pixel phase cross-correlation method. For each source, the first epoch was selected as the reference frame, and relative positional shifts between the reference image and subsequent epochs were measured using a phase cross-correlation algorithm implemented in the scikit-image555https://scikit-image.org/docs/0.25.x/api/skimage.registration.html package. An upsampling factor of 100 was adopted, corresponding to a positional accuracy of 0.01 pixel. The measured shifts were then applied via spline interpolation to align all images to the common reference frame. The results are summarized in Table B1, which lists the offsets and (in pixel units) relative to the reference epoch. Difference maps were subsequently inspected to confirm that source structures were well matched and that no systematic offsets remained.
@@
Figure 3 illustrates our source extraction procedure, taking I15520–5234 as an example. To avoid elevated edge noise introduced by primary beam correction, structure identification was performed on the continuum images uncorrected for the primary beam. Structures were extracted independently for the MaMMOtH and QUARKS datasets. For each survey, multiple observations of the same source were combined to improve sensitivity. We used the astrodendro 666http://www.dendrograms.org/ algorithm to decompose the emission hierarchically. The highest hierarchical level is a “leaf” (i.e., a structure with no substructure), corresponding to what we define as a condensation. The three key parameters are min_value (the minimum pixel intensity to be considered), min_delta (the minimum height for any local maximum to be defined as an independent entity), and min_npix (the minimum number of pixels for a leaf to be defined as an independent entity). To ensure consistency across the two surveys, we adopted a uniform set of parameters: min_value = , min_delta = –, and min_npix equals the number of pixels corresponding to the beam area. Here, denotes the rms noise level of the continuum image. For multi-epoch sources, we averaged the images prior to structure identification and determined from the averaged image before primary-beam correction. In the example shown in Figure 3, we used min_delta = , with for MaMMOtH and for QUARKS, and min_npix = 119.
@@
In Figure 3, structures extracted from the MaMMOtH and QUARKS datasets are shown with red and green contours. Although the condensations detected in the two surveys do not always match, all structures from both datasets are retained in the final union mask, shown in yellow contour. We constructed this final mask by taking the union of the individual masks of identified regions in different epochs. This union mask was then applied to each epoch for flux extraction. To ensure that the measured fluxes represent the true emission distribution, this step was performed after primary beam correction. We identified a total of 383 condensations within the 22 protoclusters. The peak intensities and coordinates of all condensations are summarized in Appendix C, where the source names indicate the corresponding region (i.e., the massive protocluster) in which each condensation is located. The peak intensity is the peak value within the dendrogram contour, which is a good approximation of the total integrated intensity for point sources. The complete table is available online. For each protocluster, condensation IDs are assigned starting from 1 and are ordered according to their spatial distribution in the map images, following a right-to-left and bottom-to-top sequence.
IV.2 Relative Calibration
@@
After aligning the observations and extracting the sources, we derived and applied a relative flux calibration factor to each dataset in order to accurately track the peak intensity variations of a given object across all epochs. This approach allows for more robust measurements of intrinsic variability within each field. Our procedure generally follows the method outlined by Mairs et al. (2017). The relative flux calibration procedure consists of the following four steps:
@@
We first divided the 22 protoclusters into six groups based on their observation dates. Protoclusters were assigned to the same group only if they were observed on exactly the same set of dates, and each group shared the same set of ALMA calibrator sources. As a result, a group may contain multiple protoclusters, while each protocluster belongs to only one group. This grouping ensures consistent temporal sampling and enables consistent relative flux calibration across all epochs within each group.
@@
At different observing epochs, even for the same field or source, the measured continuum flux may vary due to non-astrophysical effects such as changes in instrumental response, atmospheric conditions, and calibration uncertainties. To mitigate these systematic effects, we identify a subset of sources that can serve as relative flux calibrators. These sources are assumed to have intrinsically constant (or nearly constant) emission over the timescales probed by our observations. Specifically, within each group, we selected candidate calibrator sources with in all available epochs, ensuring that their measured peak intensities are robust against noise fluctuations. For each candidate, we then calculated the mean peak intensity over the valid epochs and normalized the peak intensity of each epoch by this mean. We adopted a threshold of 0.1 for this normalized standard deviation as a compromise between retaining a sufficient number of calibrator candidates and ensuring calibration reliability. Sources with normalized standard deviations below this threshold were classified as stable and used as relative flux calibrators.
@@
We then construct a normalized peak intensity ratio defined as:
(1) where is the peak intensity of source measured within the dendrogram contour at epoch , and denotes the mean over time of the peak intensity of that source, averaged over all valid epochs in the group. The flux calibration factor for each epoch was then obtained by averaging the mean flux normalized ratios across all stable calibrators in the group:
(2) This procedure assumes that, although individual calibrator sources may exhibit small residual fluctuations, their ensemble-averaged flux remains constant over time, allowing robust estimation of systematic calibration offsets. We note, however, with a limited number of calibrator sources in each group, the uncertainty on the ensemble mean may still represent a significant source of error.
@@
IV.3 Searching for (Sub)Millimeter Variables
Investigating flux variability in our survey is confronted by several key challenges. Each ALMA target was observed at only a small number of epochs (2–4 per protocluster), and the background rms varies across different protoclusters. To address these limitations, we employed two complementary methods to cross-validate variability among our samples (383 condensations). First, we adopted the methodology outlined by Johnstone et al. (2018), originally developed for the JCMT Transient Survey. Although that survey benefits from more extensive temporal coverage and highly uniform observing conditions, which result in greater statistical robustness, our use of consistent data reduction and relative flux calibration still allows us to adopt the same framework as a practical indicator for identifying candidate variable sources. Second, to further increase the reliability of our results, we applied a difference map analysis to independently verify variability in the identified candidates.
IV.3.1 Standard Deviation Analysis
For each detected condensation, we extract the peak intensity measured in each epoch and compute two key statistical quantities: the mean of the peak intensity over time, , and the corresponding standard deviation (). The uncertainty in measuring the peak intensity of these sources is primarily influenced by two factors: the background rms for faint condensations, relative calibration accuracy for bright condensations. We defined the fiducial standard deviation for each source, , which represents the expected measurement uncertainty in peak intensity, as:
where denotes the background rms associated with the source . The term represents the expected relative flux calibration uncertainty, which we adopt to be 0.5% (see Sect. IV.2 for details), while denotes the time-averaged peak intensity of the source.
Figure 5 shows the normalized standard deviation () as a function of the mean peak intensity for each condensation. In this work, we adopt a conservative variability threshold of , beyond which deviations are unlikely to result from calibration uncertainties or stochastic noise. The JCMT Transient Survey employed a lower threshold of 2 (Johnstone et al., 2018; Chen et al., 2025); however, given our smaller number of epochs, we adopted a more stringent criterion to minimize false positives and ensure that only the most statistically significant outliers are flagged as candidate variables.
| Source Name | Condensation IDa | SD/SDfid | Fractional Difference (%)b | Fractional Amplitude (%) | Typec | d | |||
| Epoch 2-1 | Epoch 3-1 | Epoch 4-1 | (mJy beam-1) | (kpc) | |||||
| I11332-6258 | 7 | 7.4 | 11.7 | 14.4 | 25.55 | 13.3 | hot core | 1.40 | |
| I12320-6122 | 5 | 6.5 | 16.2 | 17.97 | 15.0 | young protostar | 4.17 | ||
| I13111-6228 | 10 | 28.7 | 68.3 | 19.51 | 51.1 | HC Hii region | 2.97 | ||
| I16076-5134 | 6 | 5.7 | -0.8 | 7.8 | 5.3 | 65.33 | 8.3 | hot core | 5.31 |
| I16484-4603 | 1 | 7.4 | -7.3 | -9.3 | -10.4 | 97.37 | 11.1 | hot core | 2.17 |
@@ IV.3.1 Standard Deviation Analysis
Applying this criterion, we identify five candidate variables among the 383 detections with multi-epoch measurements. Their quantitative variability metrics are summarized in Table 1.
The fractional difference is calculated as , where and are the peak intensities at two different epochs. Multiple pairs of epochs (e.g., 2–1, 3–1, and 4–1) are used to evaluate the variability across the monitoring period. The observed fractional amplitude is defined as , where and denote the maximum and minimum peak intensities measured across all observing epochs, and is the mean peak intensity over time.
Among them, condensation 10 in I13111–6228 is particularly notable, exhibiting a normalized standard deviation of , far above the adopted threshold. This is the first detection of millimeter variability toward this protocluster. The other four candidates show more moderate excess deviations, with ranging from 5.7 to 7.4.
Although the fiducial standard deviation analysis provides a global identification of candidate variables, it does not explicitly capture the epoch-to-epoch flux behavior of individual condensations. To complement this approach, we extend the same fiducial standard deviation model by propagating uncertainties and computing the peak intensity ratio between epochs, referenced to epoch 1 for consistency (see Appendix D). We find that the variable candidates identified with this ratio-based method are fully consistent with those selected from the fiducial standard deviation analysis.
As an illustration, Figure 6 shows the source I13111–6228, in which condensation 10 exhibits the most significant intensity variation among all candidates. This condensation is clearly offset from the fiducial model, well exceeding the threshold, and is marked with a red star. For reference, the points with are colored green, reflecting their larger noise-dominated uncertainties, while those with higher S/N are shown in blue and provide more reliable measurements. Additional ratio plots for representative candidates are shown in Appendix D, and the complete set of figures is available online.
IV.3.2 Difference Maps Analysis